determine whether the graphs of 2 lines are parrallel
Answer by Alan3354(69443)

Find the slope of both lines by putting them in slope-intercept form,
y = mx+b.
(That means solve for y)
y+5=7x
y = 7x - 5
slope = +7
----------
and 2x-y=-5
y = 2x + 5
slope = +2
The slopes are not the same, so they are not parallel.